An observer travelling with a constant velocity of 20ms-1, passes close to a stationary source of sound and notices that there is a change of frequency of 50 Hz as he passes the source. What is the frequency of the source? 


1
Expert's answer
2021-03-14T19:47:13-0400

The source is stationary so "v_s =0" while the observer/listener's velocity is "v_L=20\ m/s" .

What is given is the change in frequency which is 50 Hz, this means "f_L=f_s+50\ Hz." Thus, from

"f_L={f_s(v+v_L) \over (v + v_s)}" we obtain


"f_S + 50Hz = \dfrac{v+v_L}{v}f_S\ \ \ \Rightarrow\ \ \ 50Hz= (\dfrac{v+v_L}{v}-1)f_S,\\f_S =(50Hz)(\dfrac{v+v_L}{v}-1)^{-1}=\ \ \ (50Hz)(\dfrac{340m/s+20m/s}{340m/s}-1)^{-1}=850\ Hz"

Therefore, the frequency of the sound emitted by the source is 850 Hz.